help me solve this fluids problem. i need an answer for what the new height would be if pressure is increased by dp.

Find new differential height if Pa is increased by dp

Pa id atmospheric pressure.
Answer #1

. Apely pressure equilibrim at line day Pi + V (atti) = TagH +Vx +Pz P-P2 = Vug H + Valk -(((x+H)) P-P2 = Fag H + x - KR+TH d